OpenStudy (anonymous):

Expand using properties of logs: ln X^2(X^2-1)/√X+1

zepdrix (zepdrix):

All of that stuff is inside of one log?

OpenStudy (anonymous):

Its a quotient problem. The x^2(x^2-1) in the top part of the quotient and the √x-1 is under

zepdrix (zepdrix):

This one is a bit of a doozy, lot of little steps.

zepdrix (zepdrix):

A rule of logs states: log(a/b) = log(a) - log(b) So ln x^2(x^2-1)/√x+1 becomes, ln x^2(x^2-1) - ln√x+1

zepdrix (zepdrix):

Understand how I applied that step? ~Separating the numerator and denominator into different logs.

OpenStudy (anonymous):

yes i'm there

zepdrix (zepdrix):

Another important rule of logs: log(b*c) = log(b) + log(c) When we multiply we can split them up into the sum of logs. Understand how we might apply this rule? :o

OpenStudy (anonymous):

how are you applying it

zepdrix (zepdrix):

In the first log, we have: ln x^2*(x^2-1) See the multiplication? we can rewrite this as 2 logs being added together.

OpenStudy (anonymous):

so x^2+x^2+x^2-1?

zepdrix (zepdrix):

Hmm I'm not sure what you did there.. we don't want to multiply it out. We want to apply the log rule: ln x^2(x^2-1) - ln√x+1 = ln x^2+ ln(x^2-1) - ln√x+1

zepdrix (zepdrix):

There's a bunch more little steps from here.

OpenStudy (anonymous):

i understand what you did where do you go from here?

OpenStudy (anonymous):

my next step was 2lnx+2lnx-1-1/2lnx+1

zepdrix (zepdrix):

Your first term looks good. 2lnx Your last term also looks good. (1/2)ln(x+1) You rewrote the root as a 1/2 power, and pulled it outside. Let's fix up the middle term.

zepdrix (zepdrix):

ln(x^2-1) We have the `difference of squares` inside of the log. We can factor the difference of squares into conjugates. (x^2-1) = (x-1)(x+1) So that gives us: ln(x^2-1) = ln(x-1)(x+1) All of that inside of the log.

zepdrix (zepdrix):

Then you might notice that again, we have multiplication. So we can write this as the sum of logs: ln(x-1) + ln(x+1) That would be our middle term.

OpenStudy (anonymous):

ok so like foil method.i got i jeeez!

OpenStudy (anonymous):

*it